# Asymptotically Euclidean Solutions of the Constraint Equations with Prescribed Asymptotics

### v2

@@ 1 Introduction
We define these fall-off conditions as type (CK). This conclusion was later generalized by Bieri lydia1 ; lydia2 to initial data sets with the relaxed fall-off conditions (type (B))
@@
In each case, one has different degrees of control on the induced fall-off rates of various curvature components as one approaches $\mathscr{I}^{\pm}$. As also investigated by Bieri lydia4 ; lydia14 ; lydia5 , different control still is afforded by the intermediate fall-off conditions (type (A))
@@
In this work, we significantly generalize the results of bieri_brill_2025 both to do away with any symmetry assumptions and to handle type (A) data, establishing a sequence of results guaranteeing that one can control the fall-off rates and the ADM momenta of initial data sets generated via the widely utilized conformal method. We use these results to construct an initial data set of type (A) whose evolution cannot satisfy the antipodal conjecture of strominger , restricting the class of spacetimes to which this conjecture could apply. Since the conjecture holds trivially for spacetimes evolved from type (CK) data, our results indicate that the domain on which the conjecture might hold nontrivial content must be somewhat small. For simplicity of presentation, we work with $\Sigma=\mathbb{R}^{n}$, but the analytical results generalize straightforwardly to all asymptotically Euclidean manifolds, even with multiple ends. Of course, one takes $n=3$ in standard general relativity (as above), but we have found it interesting and instructive to generalize the mathematical results presented here to arbitrary dimension $n\geq 3$.
We remark that Dain and Friedrich published a work in 2001 dain2001asymptotically having a fair bit of conceptual overlap with (and a remarkably similar title to) the present work. They also proved a number of interesting results allowing control of asymptotic features of initial data sets constructed via the conformal method in $n=3$, including the prescription of a leading-order Euclidean transverse traceless tensor in $\widetilde{K}_{ij}$ allowing a priori control of the ADM momentum (cf. our Theorem 5.3). However, there is little technical overlap with the present work, regarding either precisely what is proved or the analytical techniques employed. In particular, Dain and Friedrich worked (appropriately for their objectives) in the restricted setting of initial data sets admitting a conformal compactification at spatial infinity, so they were interested in obtaining asymptotics which amount to analyticity at spatial infinity, asking that $\tilde{g}_{ij}$ and $\widetilde{K}_{ij}$ admit a full power series expansion in integer powers of $1/r$. The present work seeks to provide a simple procedure for constructing initial data sets in a much broader category, with an emphasis on controlling the structure of the leading-order terms (especially in $\tilde{g}_{ij}$) and the precise (non-integral) decay rate of subleading terms, and making no demand for a full power series expansion. Broadly, we are interested in the construction of initial data sets conforming to varying notions of asymptotic flatness for the ultimate purpose of probing the extent to which physically interesting features of spacetime depend on this notion, while the results in dain2001asymptotically are not amenable to handling a variety of fall-off behaviors. The results in dain2001asymptotically could be described as allowing the construction of highly specialized examples of type (CK) data, but they are not relevant to the construction of type (A) data sets which are not (CK), which is among our chief concerns.
@@ 2 Preliminaries and The Conformal Method
We are interested in prescribing the fall-off rates of $\tilde{g}$ and $\widetilde{K}$ by means of building them into the seed data $g$, $A_{ab}$, and $\delta N$. To leverage analytical control over the decay rates of these quantities, we operate in weighted Sobolev spaces, which we now define:
@@ 5 Constructions
In attempting to repeat this procedure for either type (CK) or type (A) data, equations (1.1)-(1.2) and (1.5)-(1.6), with all the fall-off weights shifted down by one, one finds that $Z_{a}\in W^{k-2,p}_{-1/2-n}$, so that the desired domain of $P_{g,N}$ now has $\delta=3/2-n<2-n$, and we cannot simply invoke Proposition 3.1(e). This obstruction is not surprising: Proposition 3.1 indicates that we should expect the solution $W^{a}$ to equation (3.1) to be contained in $W^{k,p}_{(2-n)^{+}}$, ultimately corresponding to $\widetilde{K}_{ab}\in W^{k-1,p}_{(1-n)^{+}}$ and allowing an $O(r^{1-n})$ term in $\widetilde{K}_{ab}$ which gives rise to the initial data set’s linear ADM momentum. The same obstruction arises in seeking $\varphi$: The conclusion of Theorem 4.1 is limited due to the generic appearance of an $O(r^{2-n})$ term in $\varphi$, corresponding to the mass term in the metric and screening out control of any remaining decay. While the problem of extracting the $O(r^{2-n})$ term from $\varphi$ to control the remaining decay has been treated in the literature (see Proposition 1 of bieri_brill_2025 or Theorem 1.17 of bartnik1 ), previously existing results are insufficient to guarantee the emergence of type (A) data in which the mass term $h_{ij}$ in the metric is anisotropic, or to fully understand the $O(r^{1-n})$ term in the second fundamental form. As demonstrated in Theorems 5.2 and 5.3 below, we are nevertheless able to treat both of these concerns.
@@
$\mathcal{P}_{j}(\widetilde{K}_{ab})=\mathcal{G}_{j}\left(\frac{1}{2N}A_{ab}\right)=-\mathcal{R}_{j}\left(\frac{1}{2N}A_{ab}\right):=-\int_{\mathbb{R}^{n}}\nabla^{b}\left(\frac{1}{2N}A_{ab}\right)k^{a}_{(j)}\;dV_{g},$ (5.1)
@@
$\widetilde{K}_{ij}$ $=\frac{n(n-2)C_{n}}{r^{n-1}}\left[p_{i}n_{j}+p_{j}n_{i}+((n-2)n_{i}n_{j}-\delta_{ij})p_{k}n^{k}\right]+o_{\ell}(r^{\tau-1}).$

### v3

@@ 1 Introduction
We define these falloff conditions as type (CK). This conclusion was later generalized by Bieri lydia1 ; lydia2 to initial data sets with the relaxed falloff conditions (type (B))
@@
In each case, one has different degrees of control on the induced falloff rates of various curvature components as one approaches $\mathscr{I}^{\pm}$. As also investigated by Bieri lydia4 ; lydia14 ; lydia5 , different control still is afforded by the intermediate falloff conditions (type (A))
@@
In this work, we significantly generalize the results of bieri_brill_2025 both to do away with any symmetry assumptions and to handle type (A) data, establishing a sequence of results guaranteeing that one can control the falloff rates and the ADM momenta of initial data sets generated via the widely utilized conformal method. We use these results to construct an initial data set of type (A) whose evolution cannot satisfy the antipodal conjecture of strominger , restricting the class of spacetimes to which this conjecture could apply. Since the conjecture holds trivially for spacetimes evolved from type (CK) data, our results indicate that the domain on which the conjecture might hold nontrivial content must be somewhat small. For simplicity of presentation, we work with $\Sigma=\mathbb{R}^{n}$, but the analytical results generalize straightforwardly to all asymptotically Euclidean manifolds, even with multiple ends. Of course, one takes $n=3$ in standard general relativity (as above), but we have found it interesting and instructive to generalize the mathematical results presented here to arbitrary dimension $n\geq 3$.
@@
We remark that Dain and Friedrich published a work in 2001 dain2001asymptotically having a fair bit of conceptual overlap with (and a remarkably similar title to) the present work. They also proved a number of interesting results allowing control of asymptotic features of initial data sets constructed via the conformal method, including the prescription of a leading-order Euclidean transverse traceless tensor in $\widetilde{K}_{ij}$ allowing control of the ADM momentum (cf. our Theorem 5.3). However, there is little technical overlap with the present work, either in precisely what is proved or in the analytical techniques employed. Namely, Dain and Friedrich worked (appropriately to their objectives) in the restricted setting of initial data sets admitting a conformal compactification at spatial infinity, so they were interested in obtaining asymptotics which amount to analyticity at spatial infinity, asking that $\tilde{g}_{ij}$ and $\widetilde{K}_{ij}$ admit a full power series expansion in integer powers of $1/r$. The present work seeks to provide a simple procedure for constructing initial data sets in a much broader category, with emphasis on controlling the structure of the leading-order terms (especially in $\tilde{g}_{ij}$) and the precise (non-integral) decay rate of subleading terms, and making no demand of a full power series expansion. Broadly, we are interested in the construction of initial data sets conforming to varying notions of asymptotic flatness for the ultimate purpose of probing the extent to which physically interesting features of spacetime depend on this notion, but the results in dain2001asymptotically are not amenable to handling a variety of falloff behaviors. In particular, results in dain2001asymptotically could be described as allowing the construction of highly specialized examples of type (CK) data, but they are not relevant to the construction of type (A) data which is not (CK), which is among our chief concerns.
In a similar vein, in 1996 Beig and O’Murchadha used spherical harmonic expansions to investigate the asymptotics of solutions to the CMC vacuum momentum constraint equation (2.4) beig_momentum_1996 . Working in the smooth setting in dimension $n=3$, they fully characterized the asymptotic forms and decay rates of solutions to the momentum constraint with seed data having given arbitrarily fast, non-exceptional falloff compatible with type (CK) data. This characterization was made up to a collection of constant “multipole moments”, including the ADM momentum components, which can be prescribed via an appropriate choice of seed data (but not anticipated for an arbitrary choice of seed data). We have built upon Beig and O’Murchadha’s work by doing the following: extending some of their conclusions to lower regularity and arbitrary dimension; establishing that similar conclusions hold for solutions to the Hamiltonian constraint equation (2.3) (and hence one’s entire initial data set); showing that the ADM momentum components can be computed directly from the seed data; and investigating the asymptotics of solutions if the seed data has the exceptional “threshold” falloff $O(r^{2-n})$, allowing the construction of type (A) data which is not (CK).
@@ 2 Preliminaries and The Conformal Method
We are interested in prescribing the falloff rates of $\tilde{g}$ and $\widetilde{K}$ by means of building them into the seed data $g$, $A_{ab}$, and $\delta N$. To leverage analytical control over the decay rates of these quantities, we operate in weighted Sobolev spaces, which we now define:
@@ 5 Constructions
In attempting to repeat this procedure for either type (CK) or type (A) data, equations (1.1)-(1.2) and (1.5)-(1.6), with all the falloff weights shifted down by one, one finds that $Z_{a}\in W^{k-2,p}_{-1/2-n}$, so that the desired domain of $P_{g,N}$ now has $\delta=3/2-n<2-n$, and we cannot simply invoke Proposition 3.1(e). This obstruction is not surprising: Proposition 3.1 indicates that we should expect the solution $W^{a}$ to equation (3.1) to be contained in $W^{k,p}_{(2-n)^{+}}$, ultimately corresponding to $\widetilde{K}_{ab}\in W^{k-1,p}_{(1-n)^{+}}$ and allowing an $O(r^{1-n})$ term in $\widetilde{K}_{ab}$ which gives rise to the initial data set’s linear ADM momentum. The same obstruction arises in seeking $\varphi$: The conclusion of Theorem 4.1 is limited due to the generic appearance of an $O(r^{2-n})$ term in $\varphi$, corresponding to the mass term in the metric and screening out control of any remaining decay. While the problem of extracting the $O(r^{2-n})$ term from $\varphi$ to control the remaining decay has been treated in the literature (see Proposition 1 of bieri_brill_2025 or Theorem 1.17 of bartnik1 ), previously existing results are insufficient to guarantee the emergence of type (A) data in which the mass term $h_{ij}$ in the metric is anisotropic, or to fully understand the $O(r^{1-n})$ term in the second fundamental form. As demonstrated in Theorems 5.2 and 5.3 below, we are nevertheless able to treat both of these concerns.
@@
$p_{j}:=\mathcal{P}_{j}(\widetilde{K}_{ab})=\mathcal{G}_{j}\left(\frac{1}{2N}A_{ab}\right)=-\mathcal{R}_{j}\left(\frac{1}{2N}A_{ab}\right):=-\int_{\mathbb{R}^{n}}\nabla^{b}\left(\frac{1}{2N}A_{ab}\right)k^{a}_{(j)}\;dV_{g},$ (5.1)
@@
$\widetilde{K}_{ij}$ $=\frac{2n(n-2)C_{n}}{r^{n-1}}\left[p_{i}n_{j}+p_{j}n_{i}+((n-2)n_{i}n_{j}-\delta_{ij})p_{k}n^{k}\right]+o_{\ell}(r^{\tau-1}).$